OpenStudy (anonymous):

(2radical7 + radical 11) (radical7-3 radical 11) multiply and simplify. assume all variables represent nonnegative numbers

OpenStudy (anonymous):

\[\frac{2\sqrt{7}+\sqrt{11}}{\sqrt{7}-3\sqrt{11}}\]?

OpenStudy (anonymous):

if so multiply top and bottom by "conjugate" of denominator, namely \[\sqrt{7}+3\sqrt{11}\]

OpenStudy (anonymous):

because \[(\sqrt{7}-3\sqrt{11})(\sqrt{7}+3\sqrt{11})=7-9\times 11=7-99=-92\]

OpenStudy (anonymous):

you get \[\frac{2\sqrt{7}+\sqrt{11}}{\sqrt{7}-3\sqrt{11}}\times \frac{\sqrt{7}+3\sqrt{11}}{\sqrt{7}+3\sqrt{11}}\]?

OpenStudy (anonymous):

\[=\frac{14+7\sqrt{77}+33}{-92}\] \[=\frac{47+7\sqrt{7}}{-92}\]

OpenStudy (anonymous):

nope not right.